# Accelerated Test-Time Scaling with Model-Free Speculative Sampling

### v1

@@ Abstract
Language models have demonstrated remarkable capabilities in reasoning tasks through test-time scaling techniques like best-of-N sampling and tree search. However, these approaches often demand substantial computational resources, creating a critical trade-off between performance and efficiency. We introduce STAND (STochastic Adaptive N-gram Drafting), a novel model-free speculative decoding approach that leverages the inherent redundancy in reasoning trajectories to achieve significant acceleration without compromising accuracy. Our analysis reveals that reasoning paths frequently reuse similar reasoning patterns, enabling efficient model-free token prediction without requiring separate draft models. By introducing stochastic drafting and preserving probabilistic information through a memory-efficient logit-based N-gram module, combined with optimized Gumbel-Top-K sampling and data-driven tree construction, STAND significantly improves token acceptance rates. Extensive evaluations across multiple models and reasoning tasks (AIME-2024, GPQA-Diamond, and LiveCodeBench) demonstrate that STAND reduces inference latency by 60-65% compared to standard autoregressive decoding while maintaining accuracy. Furthermore, STAND outperforms state-of-the-art speculative decoding methods by 14-28% in throughput and shows strong performance even in single-trajectory scenarios, reducing inference latency by 48-58%. As a model-free approach, STAND can be applied to any existing language model without additional training, being a powerful plug-and-play solution for accelerating language model reasoning.
@@
Woomin Song1,2,†,   Saket Dingliwal1,   Sai Muralidhar Jayanthi1, Bhavana Ganesh1, Jinwoo Shin2, Aram Galstyan1, Sravan Babu Bodapati1 1Amazon AGI 2KAIST
00footnotetext: Work done during an internship at Amazon.
@@ 1 Introduction
Test-time scaling has emerged as a prominent paradigm for enhancing the performance of language models by allocating additional computational resources during inference (Snell et al., 2024). This includes generating long sequences of thoughts though Large Reasoning Models (LRMs) Muennighoff et al. (2025), multi-sampling approaches like best-of-N sampling and majority voting that generate multiple independent outputs to select the most promising one (Wang et al., 2022), as well as iterative methods like tree search and sequential refinement that allow models to progressively improve their reasoning process (Uesato et al., 2022). While these methods demonstrate the potential for significant accuracy improvements through increased computation, they often demand substantial computational resources due to the large number of tokens that need to be generated.
Recent research has focused on reducing the high computational costs of test-time scaling and reasoning approaches (Sui et al., 2025). Some work has explored training with length-based rewards to generate more concise outputs (Aggarwal and Welleck, 2025; Qu et al., 2025), while other approaches use combinations of small and large models to distribute the workload efficiently (Liao et al., 2025; Yang et al., 2025). However, these efficiency-focused methods typically face a fundamental trade-off. While they reduce computational costs, they tend to sacrifice some accuracy compared to more exhaustive approaches, as using fewer samples or cutting short the exploration process often leads to lower performance.
@@
This raises a crucial question: How can we improve the efficiency of test-time scaling and reasoning methods without compromising their accuracy? To address this challenge, we turn our attention to speculative decoding (SD), which offers a promising solution for lossless acceleration of language model inference. Speculative decoding accelerates language model inference by using a smaller "draft" model to predict tokens, which are then verified by the larger target model (Leviathan et al., 2023). With appropriate verification strategies (Chen et al., 2023a), SD can speed up the auto-regressive decoding process of large language models while preserving their output distribution.
A key observation in LRMs is the significant repetition of token sequences across different reasoning paths. When models are performing chain-of-thought reasoning (Snell et al., 2024) or exploring multiple solutions (Wang et al., 2022; Xie et al., 2024), they frequently reuse similar expressions, logical deductions, and reasoning patterns.
This redundancy presents an opportunity for model-free speculative decoding (Saxena, 2023; Ou et al., 2024). Unlike model-based approaches that rely on neural networks as drafters (Li et al., 2024c; Cai et al., 2024), model-free methods can leverage patterns from previous generations to construct drafts. This makes them particularly well-suited for exploiting cross-trajectory information. Our experiments confirm this approach’s effectiveness, demonstrating improved efficiency as the number of reasoning trajectories increases.
@@
Our extensive evaluations demonstrate STAND’s effectiveness in various reasoning tasks (math, science and coding) and different model scales. As highlighted in Figure 1, our results show that STAND’s benefits become more pronounced as the number of reasoning trajectories increases. When using best-of-16 sampling to achieve optimal accuracy, STAND reduces inference latency by 60-65% compared to standard autoregressive decoding while maintaining the same performance level. Moreover, STAND outperforms state-of-the-art speculative decoding methods by 14-28% in throughput, establishing an efficient drafting strategy for reasoning tasks.
Furthermore, we observe that STAND also shows the best throughput in single-trajectory evaluations, reducing the inference latency by 48-58% compared to standard autoregressive decoding, although it was primarily designed to leverage information across multiple reasoning trajectories. As a model-free speculative decoding approach, STAND accomplishes all these achievements without requiring any additional drafter model, or fine-tuning the target model, being able to be used in plug-and-play manner to any existing LRMs.
@@ 2 Related Works
Test-Time Scaling (TTS) has emerged as a prominent strategy to enhance problem-solving capabilities during inference without model retraining Snell et al. (2024); Muennighoff et al. (2025); Wang et al. (2022); Uesato et al. (2022); Xie et al. (2024). Generating long chain-of-thoughts or sampling multiple sequences have consistently showcased higher accuracy in many complex tasks like math, science and coding Wei et al. (2022); Cobbe et al. (2021); Chen et al. (2023b). However, the computational cost of TTS remains a critical bottleneck for their practical use. Recent work has explored optimizing inference using adaptive thinking lengths, cascading models of different sizes, length penalties during training, and budget-constrained decoding Aggarwal and Welleck (2025); Qu et al. (2025); Liao et al. (2025); Li et al. (2024b); Wan et al. (2024), yet the fundamental trade-off between accuracy gains and costs persists. Our method aims to accelerate reasoning while ensuring no performance degradation.
@@
SD have been shown to accelerate Large Language Model (LLM) inference without any loss in the accuracy Leviathan et al. (2023). The approach typically involves a smaller "draft" model proposing candidate token sequences for parallel verification by the larger "target" model. If the tokens align with the target model’s output distribution, they are "accepted", resulting in more than one token being produced in a single forward pass of the LLM. Various compute-efficient drafting strategies have been proposed in the literature to increase the chances of acceptance. Neural draft architectures have evolved from simple, smaller LMs to sophisticated self-drafting approaches e.g., Medusa Cai et al. (2024), Eagle Li et al. (2024d), and ReDrafter Cheng et al. (2024)). Although the use of light-weight model-free drafters based on n-grams Li et al. (2024a); Somasundaram et al. (2024); Hu et al. (2024); Oliaro et al. (2024); Geva et al. (2023); Ou et al. (2024); Saxena (2023) has been explored previously for generic tasks, we revisit them in the context of LRMs. While these approaches limit themselves to deterministic n-gram based lookups as draft sequences, we highlight the significance of stochastic drafting with logit information of previously generated n-grams for reasoning in our proposed method. To further boost SD performance, tree drafting was proposed where multiple draft token predictions are organized in a tree structure, enabling efficient parallel verification through a specialized tree attention mask Miao et al. (2023); Li et al. (2024d). Methods like Eagle-2 Li et al. (2024c) even used dynamic tree layout choices for SD. Extending these existing methods, we additionally propose a computationally efficient data-driven offline tree optimization method for our lightweight model-free drafting method for LRMs.
Other approaches in literature that tie SD with LRMs include Speculative Thinking Hu et al. (2025), SpecReason Pan et al. (2025), Reward-guided SD Liao et al. (2025). However, they do not maintain the lossless nature of SD and hence can also be used in combination with our work. A contemporary work Li et al. (2025) have explored the importance of model-free n-gram based drafting for multi-sample inference, but did not showcase any practical speedup. We extend their findings with our novel model-free stochastic drafting, and showcase a comparative analysis with existing methods through our extensive experimentation.
@@ 3.1 N-gram overlap analysis
To assess the degree of redundancy in reasoning trajectories, we conducted a comprehensive analysis of n-gram overlap patterns across multiple solutions generated by the DeepSeek-R1-Distill-Qwen-7B model on the AIME-2024 dataset. Figure 2 illustrates our findings, depicting the overlap rates for n-grams ranging from bigrams to 5-grams across varying numbers of reasoning trajectories.
@@ 3.2 Effectiveness of stochastic drafting
In contrast to traditional generation approaches that rely on greedy decoding, LRMs typically employ sampling-based generation strategies to produce multiple diverse solution trajectories, making the choice of drafting strategy particularly crucial. In speculative sampling Chen et al. (2023a), given a target distribution $p(x)$ and draft distribution $q(x)$, the speculative sampling procedure operates by first sampling $x\sim q(x)$. The sampled token is accepted if $q(x)\leq p(x)$. Otherwise, when $q(x)>p(x)$, the token is rejected with probability $1-\frac{p(x)}{q(x)}$ and resampled from an adjusted distribution $p^{\prime}(x)=\text{norm}(\max(0,p(x)-q(x)))$. This procedure guarantees that the final output distribution matches the target distribution $p(x)$, for any drafting distribution $q(x)$.
One can choose the drafting strategy to be deterministic or stochastic. In the former, $q(x)$ is treated as a one-hot vector where $q(x_{\text{draft}})=1$ for the drafted token $x_{\text{draft}}$ and $q(x)=0$ for all other $x$. For speculative sampling, this means the drafted token is accepted with $p(x_{\text{draft}})$, which can be particularly low when the target model is uncertain about its prediction. In contrast, stochastic drafting generates drafts through sampling from a probability distribution. Aligning this draft distribution with the target can significantly boost the chances of acceptance.
In generic greedy decoding setups where this choice does not matter, existing model-free SD methods Ou et al. (2024); Hu et al. (2024); Saxena (2023) do not store any probability distribution with the n-gram lookup-based drafters. Eagle-2 Li et al. (2024c) also uses deterministic drafting for better compatibility with their dynamic tree construction logic. However, for LRMs where sampling plays a key role in generation, we showcase that this choice plays a pivotal role in acceptance probability of the draft sequence. As shown in Figure 3, this fundamental difference leads to 5%, 7% and 8% higher acceptance probabilities for stochastic drafting compared to deterministic drafting across different reasoning tasks i.e. AIME, GPQA and LiveCodeBench respectively. These experimental findings motivated us to find effective ways to compute draft model probabilities in STAND, that aligns well with the probability distributions of LRMs from which the multiple trajectories are sampled.
@@ 4 STAND
In this section, we present the details of STAND. In Section 4.1, we propose a memory- and compute-efficient approach to construct the logit-based N-gram module. Then in Section 4.2, we illustrate how to use the N-gram module as a drafter for stochastic sampling, together with several optimization techniques that further improve performance.
@@ 4.1 Logit-based adaptive N-gram module
Traditional N-gram modules for speculative decoding typically store pairs of N-grams and their corresponding next tokens (Ou et al., 2024). We improve this approach by instead storing the logit distribution from which the next token is sampled. This modification preserves the rich probabilistic information of potential next tokens, enabling more sophisticated stochastic drafting strategies. While existing methods like Token Recycle partially utilize logit information by storing top-k token IDs, they discard valuable probability information that are crucial for stochastic drafting. Like previous works (Saxena, 2023; Ou et al., 2024), we maintain separate lookup tables from unigrams to 4-grams.
@@ 4.2 Drafting with STAND
For each position in the draft tree, we predict the next tokens using a multi-level N-gram approach. Following previous works (Saxena, 2023; Ou et al., 2024), we search for matching N-grams in decreasing order of length, from 4-grams down to unigrams, using the first successful match. This lookup returns the top-10 candidate tokens and their corresponding probabilities from our stored distributions. Based on the number of children required at each tree node, we sample k tokens without replacement from these candidates. These sampled tokens then undergo standard speculative sampling verification to ensure draft quality.
@@
For efficient stochastic drafting, we replace traditional sequential sampling with a parallel sampling approach based on the Gumbel-Top-K trick (Kool et al., 2019). For each candidate token’s log probability $\phi_{i}$, we add Gumbel noise to create a perturbed distribution:
@@
We address this limitation through a data-driven approach to static tree construction. Our method begins by initializing a large tree with 625 nodes and performing speculative decoding on 30 samples from the AIME-2024 dataset. During this process, we track which nodes are frequently part of successful speculation paths. We then select the top-80 most effective nodes and reorganize them into a compact tree structure. This empirical approach maintains the computational efficiency of static trees while ensuring the tree structure is optimized based on real-world performance data.
@@ 5 Experiments
This section highlights the effectiveness of STAND through extensive experiments. In Section 5.1, we showcase that STAND can significantly speed up generation in multi-sampling inference. Section 5.2, we highlight that STAND can also be used in single-trajectory inference, involving a single long chain-of-thought. Finally in Section 5.3, we perform an ablation study of the components that make STAND effective, followed by an additional analysis of the optimized tree structure.
@@
For model-free baselines, we compare STAND against Prompt Lookup Decoding (PLD, Saxena (2023)), Adaptive N-gram Parallel Decoding (ANPD, Ou et al. (2024)), Token Recycle (Recycle, Luo et al. (2024)), SAM-Decoding (SAM, Hu et al. (2024)) and a combination of SAM decoding and Token Recycle, also proposed in the SAM paper. For Static SAM (which is a component of SAM that uses a pre-constructed suffix automation from a datastore), we construct the datastore using 4k samples from the OpenThoughts-114k (Team, 2025) dataset. We also compare against Eagle-2, as a representative model-based baseline. To enable long context inference, we trained all Eagle models using OpenThoughts-114k dataset, where long samples exceeding 32k tokens were truncated.
@@
We adopt throughput and acceptance length as the main evaluation metrics. Throughput is obtained by dividing the total drafting time with the number of generated tokens. Acceptance length describes how many tokens were accepted per speculation step. Higher values are better for both metrics.
5.1 Main results

### v2

@@ Abstract
Language models have demonstrated remarkable capabilities in reasoning tasks through test-time scaling techniques like best-of-N sampling and tree search. However, these approaches often demand substantial computational resources, creating a critical trade-off between performance and efficiency. We introduce STAND (STochastic Adaptive N-gram Drafting), a novel model-free speculative decoding approach that exploits the inherent redundancy in reasoning trajectories to achieve significant acceleration without compromising accuracy. Our analysis shows that reasoning paths frequently reuse similar reasoning patterns, enabling efficient model-free token prediction without requiring separate draft models. By introducing stochastic drafting and preserving probabilistic information through a memory-efficient logit-based N-gram module, combined with optimized Gumbel-Top-K sampling and data-driven tree construction, STAND significantly improves token acceptance rates. Extensive evaluations across multiple models and reasoning tasks (AIME-2024, GPQA-Diamond, and LiveCodeBench) demonstrate that STAND reduces inference latency by 60-65% compared to standard autoregressive decoding while maintaining accuracy. Furthermore, STAND consistently outperforms state-of-the-art speculative decoding methods across diverse inference patterns, including single-trajectory decoding, batch decoding, and test-time tree search. As a model-free approach, STAND can be applied to any existing language model without additional training, making it a powerful plug-and-play solution for accelerating language model reasoning.
@@
Woomin Song1,†,   Saket Dingliwal2,   Sai Muralidhar Jayanthi2, Bhavana Ganesh3,‡, Jinwoo Shin1, Aram Galstyan2, Sravan Babu Bodapati2 1KAIST 2Amazon AGI 3AirSignal
00footnotetext: Work done during an internship at Amazon. Work done at Amazon.
@@ 1 Introduction
Test-time scaling has emerged as a prominent paradigm for enhancing the performance of language models by allocating additional computational resources during inference (snell2024scaling). This includes generating long sequences of thoughts though Large Reasoning Models (LRMs) muennighoff2025s1, multi-sampling approaches like best-of-N sampling and majority voting that generate multiple independent outputs to select the most promising one (Wang2022SelfConsistencyIC), as well as iterative methods like tree search and sequential refinement that allow models to progressively improve their reasoning process (uesato2022solving). While these methods demonstrate the potential for significant accuracy improvements through increased computation, they often demand substantial computational resources due to the large number of tokens that need to be generated.
Recent research has focused on reducing the high computational costs of test-time scaling and reasoning approaches (sui2025stop). Some work has explored training with length-based rewards to generate more concise outputs (aggarwal2025l1; qu2025optimizing), while other approaches use combinations of small and large models to distribute the workload efficiently (liao2025reward; yang2025speculative). However, these efficiency-focused methods typically face a fundamental trade-off. While they reduce computational costs, they tend to sacrifice some accuracy compared to more exhaustive approaches, as using fewer samples or cutting short the exploration process often leads to lower performance.
@@
This raises a crucial question: How can we improve the efficiency of test-time scaling and reasoning methods without compromising their accuracy? To address this challenge, we turn our attention to speculative decoding (SD), which offers a promising solution for lossless acceleration of language model inference. Speculative decoding accelerates language model inference by using a smaller "draft" model to predict tokens, which are then verified by the larger target model (leviathan2023fast). With appropriate verification strategies (chen2023accelerating), SD can speed up the auto-regressive decoding process of large language models while preserving their output distribution.
A key observation in LRMs is the significant repetition of token sequences across different reasoning paths. When models are performing chain-of-thought reasoning (snell2024scaling) or exploring multiple solutions (Wang2022SelfConsistencyIC; xie2024monte), they frequently reuse similar expressions, logical deductions, and reasoning patterns.
This redundancy presents an opportunity for model-free speculative decoding (pld; anpd). Unlike model-based approaches that rely on neural networks as drafters (li2024eagle2; cai2024medusa), model-free methods can leverage patterns from previous generations to construct drafts. This makes them particularly well-suited for exploiting cross-trajectory information. Our experiments confirm this approach’s effectiveness, demonstrating improved efficiency as the number of reasoning trajectories increases.
@@
Our extensive evaluations demonstrate STAND’s effectiveness across diverse reasoning tasks (math, science, and coding) and model scales. As highlighted in Figure 1, STAND’s benefits become more pronounced as the number of reasoning trajectories increases. With best-of-16 sampling for optimal accuracy, STAND reduces inference latency by 60–65% compared to standard autoregressive decoding while maintaining performance. Moreover, STAND outperforms state-of-the-art speculative decoding methods by 14–28% in throughput, establishing an efficient drafting strategy for reasoning tasks.
Furthermore, STAND consistently achieves the best throughput across multiple inference patterns, reducing the inference latency by 58% in single-trajectory decoding, 30% in batch decoding, and 61% in test-time tree search, compared to standard autoregressive decoding. As a model-free speculative decoding approach, STAND accomplishes all these achievements without requiring any additional drafter model, or fine-tuning the target model, being able to be used in plug-and-play manner to any existing LRMs.
@@ 2 Related Works
Test-Time Scaling (TTS) has emerged as a prominent strategy to enhance problem-solving capabilities during inference without model retraining snell2024scaling; muennighoff2025s1; Wang2022SelfConsistencyIC; uesato2022solving; xie2024monte. Generating long chain-of-thoughts or sampling multiple sequences have consistently showcased higher accuracy in many complex tasks like math, science and coding wei2022chain; cobbe2021training; chen2023universal. However, the computational cost of TTS remains a critical bottleneck for their practical use. Recent work has explored optimizing inference using adaptive thinking lengths, cascading models of different sizes, length penalties during training, and budget-constrained decoding aggarwal2025l1; qu2025optimizing; liao2025reward; li2024escape; wan2024dynamic, yet the fundamental trade-off between accuracy gains and costs persists. Our method aims to accelerate reasoning while ensuring no performance degradation.
@@
SD have been shown to accelerate Large Language Model (LLM) inference without any loss in the accuracy leviathan2023fast. The approach typically involves a smaller "draft" model proposing candidate token sequences for parallel verification by the larger "target" model. If the tokens align with the target model’s output distribution, they are "accepted", resulting in more than one token being produced in a single forward pass of the LLM. Various compute-efficient drafting strategies have been proposed in the literature to increase the chances of acceptance. Neural draft architectures have evolved from simple, smaller LMs leviathan2023fast; choi2025mamba to sophisticated self-drafting approaches cai2024medusa; li2024eagle; cheng2024recurrent). Although the use of light-weight model-free drafters based on n-grams li2024nearest; somasundaram2024pld+; hu2024sam; oliaro2024suffixdecoding; geva2023token; anpd; pld has been explored previously for generic tasks, we revisit them in the context of LRMs. While these approaches limit themselves to deterministic n-gram based lookups as draft sequences, we highlight the significance of stochastic drafting with logit information of previously generated n-grams for reasoning in our proposed method. To further boost SD performance, tree drafting was proposed where multiple draft token predictions are organized in a tree structure, enabling efficient parallel verification through a specialized tree attention mask Miao2023SpecInfer; li2024eagle. Methods like Eagle-2 li2024eagle2 even used dynamic tree layout choices for SD. Extending these existing methods, we additionally propose a computationally efficient data-driven offline tree optimization method for our lightweight model-free drafting method for LRMs.
Other approaches in literature that tie SD with LRMs include Speculative Thinking hu2025speculative, SpecReason pan2025specreason, Reward-guided SD liao2025reward. However, they do not maintain the lossless nature of SD and hence can also be used in combination with our work. A contemporary work li2025speculative have explored the importance of model-free n-gram based drafting for multi-sample inference, but did not showcase any practical speedup. We extend their findings with our novel model-free stochastic drafting, and showcase a comparative analysis with existing methods through our extensive experimentation.
@@ 3.1 N-gram overlap analysis
To assess the degree of redundancy in reasoning trajectories, we conducted a comprehensive analysis of n-gram overlap patterns across multiple solutions generated by the DeepSeek-R1-Distill-Qwen-7B model on the AIME-2024 dataset. Figure˜2 illustrates our findings, depicting the overlap rates for n-grams ranging from bigrams to 5-grams across varying numbers of reasoning trajectories.
@@ 3.2 Effectiveness of stochastic drafting
In contrast to traditional generation approaches that rely on greedy decoding, LRMs typically employ sampling-based generation strategies to produce multiple diverse solution trajectories, making the choice of drafting strategy particularly crucial. In speculative sampling chen2023accelerating, given a target distribution $p(x)$ and draft distribution $q(x)$, the speculative sampling procedure operates by first sampling $x\sim q(x)$. The sampled token is accepted if $q(x)\leq p(x)$. Otherwise, when $q(x)>p(x)$, the token is rejected with probability $1-\frac{p(x)}{q(x)}$ and resampled from an adjusted distribution $p^{\prime}(x)=\text{norm}(\max(0,p(x)-q(x)))$. This procedure guarantees that the final output distribution matches the target distribution $p(x)$, for any drafting distribution $q(x)$.
One can choose the drafting strategy to be deterministic or stochastic. In the former, $q(x)$ is treated as a one-hot vector where $q(x_{\text{draft}})=1$ for the most probable token $x_{\text{draft}}$ and $q(x)=0$ for all other $x$. For speculative sampling, this means the drafted token is accepted with $p(x_{\text{draft}})$, which can be particularly low when the target model is uncertain about its prediction. In contrast, stochastic drafting generates drafts through sampling from a probability distribution. Aligning this draft distribution with the target can significantly boost the chances of acceptance.
In generic greedy decoding setups where this choice does not matter, existing model-free SD methods anpd; hu2024sam; pld do not store any probability distribution with the n-gram lookup-based drafters. Eagle-2 li2024eagle2 also uses deterministic drafting for better compatibility with their dynamic tree construction logic. However, for LRMs where sampling plays a key role in generation, we showcase that this choice plays a pivotal role in acceptance probability of the draft sequence. As shown in Figure 3, this fundamental difference leads to 5%, 7% and 8% higher acceptance probabilities for stochastic drafting compared to deterministic drafting across different reasoning tasks i.e. AIME, GPQA and LiveCodeBench respectively. These experimental findings motivated us to find effective ways to compute draft model probabilities in STAND, that aligns well with the probability distributions of LRMs from which the multiple trajectories are sampled.
@@ 4 STAND
In this section, we present the details of STAND. In Section˜4.1, we propose a memory- and compute-efficient approach to construct the logit-based N-gram module. Then in Section˜4.2, we illustrate how to use the N-gram module as a drafter for stochastic sampling, together with several optimization techniques that further improve performance.
@@ 4.1 Logit-based adaptive N-gram module
Traditional N-gram modules for speculative decoding typically store pairs of N-grams and their corresponding next tokens (anpd). We improve this approach by instead storing the logit distribution from which the next token is sampled. This modification preserves the rich probabilistic information of potential next tokens, enabling more sophisticated stochastic drafting strategies. While existing methods like Token Recycle partially utilize logit information by storing top-k token IDs, they discard valuable probability information that are crucial for stochastic drafting. Like previous works (pld; anpd), we maintain separate lookup tables from unigrams to 4-grams.
@@ 4.2 Drafting with STAND
For each position in the draft tree, we predict the next tokens using a multi-level N-gram approach. Following previous works (pld; anpd), we search for matching N-grams in decreasing order of length, from 4-grams down to unigrams, using the first successful match. This lookup returns the top-10 candidate tokens and their corresponding probabilities from our stored distributions. Based on the number of children required at each tree node, we sample k tokens without replacement from these candidates. These sampled tokens then undergo standard speculative sampling verification to ensure draft quality.
@@
For efficient stochastic drafting, we replace traditional sequential sampling with a parallel sampling approach based on the Gumbel-Top-K trick (gumbeltopk). For each candidate token’s log probability $\phi_{i}$, we add Gumbel noise to create a perturbed distribution:
@@
We address this limitation through a data-driven approach to static tree construction. Our method begins by initializing a large tree with 625 nodes and performing speculative decoding on 30 data samples. During this process, we track which nodes are frequently part of successful speculation paths. We then select the top-80 most effective nodes and reorganize them into a compact tree structure. This empirical approach maintains the computational efficiency of static trees while ensuring the tree structure is optimized based on real-world performance data.
@@ 5 Experiments
This section highlights the effectiveness of STAND through extensive experiments. In Section˜5.1, we showcase that STAND can significantly speed up generation in multi-trajectory inference. Section˜5.2, we highlight that STAND can also be used in single-trajectory inference. In Section˜5.3, we extend our evaluations to more diverse inference patterns, namely batch decoding and test-time tree search. Finally in Section˜5.4, we perform an ablation study of the components that make STAND effective, followed by an additional analysis of the optimized tree structure.
@@
For model-free baselines, we compare STAND against Prompt Lookup Decoding (PLD, pld), Adaptive N-gram Parallel Decoding (ANPD, anpd), Token Recycle (Recycle, luo2024turning), SAM-Decoding (SAM, hu2024sam) and a combination of SAM decoding and Token Recycle, also proposed in the SAM paper. For Static SAM (which is a component of SAM that uses a pre-constructed suffix automation from a datastore), we construct the datastore using 4k samples from the OpenThoughts-114k (openthoughts) dataset. For all methods involving static draft trees, we apply our tree optimization algorithm using 30 samples from the AIME-2024 dataset unless otherwise stated. We also compare against Eagle-2, as a representative model-based baseline. To enable long context inference, we trained all Eagle models using OpenThoughts-114k dataset, where long samples exceeding 32k tokens were truncated.
To ensure fair comparison, we conducted all experiments on a unified codebase, adapted from Spec-Bench (specbench). In this implementation, all components such as model forwarding and draft verification are shared, and the only differences lie in the drafting algorithms themselves. Furthermore, the draft length was fixed at 80 tokens for all methods, including our proposed approach and the baselines.
@@
We adopt throughput and acceptance length as our main evaluation metrics. Throughput measures the number of tokens generated per second, computed as the total number of generated tokens divided by the total drafting time. The acceptance length quantifies the average number of tokens generated per speculation step. For both metrics, higher values indicate better performance.
5.1 Evaluation on multi-trajectory decoding